OpenStudy (anonymous):

lim x--> infinity ((7-x^3)/(2x^3-5x))

OpenStudy (anonymous):

= (7/x^30 - 1 ----------- = 0-1 2 - (5/x) --- = -1/2 2-0

OpenStudy (anonymous):

first line : - thats (7/x^3) - 1

OpenStudy (anonymous):

is it 2x^3-5x}\]? \[\lim_{x\rightarrow \infty}\frac{7-x^3}{2x^3-5x}\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok, what jimmyrep said

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw - thnx for confirming satellite - my knowledge of limits is improving i've always found them the most difficult part of calculus

OpenStudy (anonymous):

Nice job jimmy